OpenStudy (anonymous):

I need some help solving a couple of Intermediate algebra homework questions. Can anyone help me?

OpenStudy (anonymous):

\[ find F[g(x)] for the functions: F(x)=4x+9; G(x)=-5x+7\]

OpenStudy (anonymous):

Treat the x in F(x) like a "blah". we are given F(x) = 4x + 9 so F(blah) = 4*blah + 9. doesnt matter what you put into the function for blah, like a machine. so here the "blah" is g(x). so F[ g(x)] = 4 * g(x) + 9 = 4 (-5x+7) + 9 = 4*-5x + 4*7 + 9 = -20x +37 so F(g(x)) = -20x + 37